OpenStudy (anonymous):

Write an equation in point-slope form for the line through the given point with the given slope. (8,3);m=6

OpenStudy (anonymous):

(y-y1)/(x-x1)=6 (y-3)/(x-8)=6 y-3=6x-48 y=6x-45

OpenStudy (anonymous):

y-y1=6(x-x1) should be

OpenStudy (anonymous):

never seenit divided

OpenStudy (anonymous):

y-3=6(x-8)

OpenStudy (anonymous):

y-3=6x-48

OpenStudy (anonymous):

y=6x-45

OpenStudy (anonymous):

guess it works out o the same though :oP